OpenStudy (anonymous):

Use the arc length formula and the given information to find s. r = 20 ft θ = 19° s = ?

OpenStudy (anonymous):

\[\frac{ 19 }{ 360 }=\frac{ s }{ 2*\pi*20}\]

OpenStudy (anonymous):

can you show me the work

OpenStudy (anonymous):

@marshallinwashington

OpenStudy (anonymous):

Let pi = 3.14 So to solve proportions you use either algebra or you can cross multiply and divide by the the third number. So cross multiplying, you have 19*2*3.14*20 which equals 2,386.4 Now divide by the third number, 360. 2,386.4 divided by 360 will get you 6.63 as your answer. Since the units are in feet, your answer will be 6.63 feet

OpenStudy (anonymous):

See this you tube video on solving proportions. https://www.youtube.com/watch?v=Ss3fp51ghMo